If a rock falls from a height of 20 meters on Earth, the height H (in meters) after x seconds is approximately
H(x) = 20 - 4.9x2.
When does the rock strike the ground? Round to the nearest hundredth, if necessary.
A. 4.08 sec
B. 0.91 sec
C. 0.83 sec
D. 2.02 sec


Answer by ewatrrr(24785)


Hi
H(x) = 20 - 4.9x^2 |Height H (in meters) after x seconds
When does the rock strike the ground?
0 = 20 -4.9x^2
20/4.9 = x^2
x = 2.02sec